Lets be the directed straight line such that the initial point is l-i and the terminal point is liti. Find I dz f ze - 22

complex anaylsis
Answer #1

Explanation 1 dz 2 22 P line is directed straight where, I foom 1-0 to 1+ Img (1+i) 1 Re - So, Z= (1+i) dz = Z(2-2) z=(1-0)In (2-2) filma) = ln (m) - In (o) } In (1+i) - In (1-0) -In (1+1-2) + ln (181-3) 11 = Incite) - Inc-e) - Jh (1-1) + 6:-) -> [io coso tisine - - 1/2 16-09] { 2 fon Ancie) - da (-os) In(es) - ba (214 = UTC + 2 2 it = IN - ir 2 ANSWER ... =